2. You want a new car that costs $32,000. Figure out the following: ______________ What would the monthly payment be if you borrowed $32,000 from a credit union at 4% for 5 years? ______________ What would be the total of monthly payments (the amount you would actually pay for the car)? ______________ How much interest would you pay if you did this? ______________ How much would you have to save each month if you decided to save for 5 years and pay cash for the car, and you could earn 6% on your savings? ______________ What would be the total of monthly savings (the amount you would actually pay for the car if you saved)?
Answer #1

Monthly loan payment is calculated using PMT function in Excel :

rate = 4% / 12   (converting annual rate into monthly rate)

nper = 5*12 (5 year loan with 12 monthly payments each year)

pv = 32000 (loan amount)

PMT is calculated to be $589.33

A1 - * fac : B =PMT(4%/12,5*12,32000) D E F C A ($589.33) 1

Total of monthly payments = monthly payment * number of years * number of months per year

Total of monthly payments = $589.33 * 5 * 12 = $35,359.72

Interest paid = total of monthly payments - cost of car

Interest paid = $35,359.72 - $32,000 = $3,359.72

Monthly saving is calculated using PMT function in Excel :

rate = 6% / 12   (converting annual rate into monthly rate)

nper = 5*12 (5 years of saving period with 12 monthly savings each year)

pv = 0 (amount already saved currently is zero)

fv = 32000 (cost of car)

PMT is calculated to be $458.65

A3 X fac =PMT(6%/12,5*12,0,32000) F 3) ($458.65)!

Total of monthly savings = monthly saving amount * number of years * number of months per year

Total of monthly savings = $458.65 * 5 * 12

Total of monthly savings = $27,518.98